OpenStudy (anonymous):

may someone please explain how to solve z/5+3= -35

OpenStudy (whpalmer4):

is that \[\frac{z}{5}+3=-35\]or\[\frac{z}{5+3}=-35\]

OpenStudy (whpalmer4):

ok, general rule of thumb is that you can do just about anything to an equation as long as you do it to both sides. another is to move everything you can with addition and subtraction, getting every term involving your variable on one side of the - with nothing else there, then do multiplication or division to get rid of anything in front of or underneath your variable. I would start by subtracting \(3\) from both sides. What do you get when you do that?

OpenStudy (anonymous):

i get z over 5 and 38 over 4

OpenStudy (whpalmer4):

uh, how do you get 38 over 4? -35-3=

OpenStudy (whpalmer4):

\[\frac{z}5+3-3=-35-3\]\[\frac{z}5\cancel{+3-3}=-38\]\[\frac{z}5=-38\]what is the next step?

OpenStudy (whpalmer4):

no, you need to multiply both sides of the equation by 5. what do you get when you do so?

OpenStudy (anonymous):

i get z= -190